OpenStudy (anonymous):

Which expression is a factor of 4ab + 4a – 3b – 3? b – 1 4a – 3 3b + 1 4a + 3b

OpenStudy (anonymous):

i got b

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@mathstudent55

OpenStudy (ivettef365):

when you split in two you have 4a(b+1) - 3(b+1)

OpenStudy (ivettef365):

that is equal to (4a-3)(b+1)

OpenStudy (anonymous):

so i was right?

OpenStudy (ivettef365):

so b is correct, the factors are (4a-3)(b+1)

OpenStudy (anonymous):

y varies inversely as twice x. When x = 3, y = 6. What is the value of y when x = 9? 6 3 2 9

OpenStudy (anonymous):

i got c

OpenStudy (ivettef365):

yes answer is 2

OpenStudy (mathstudent55):

Inverse variation is: y = k/x for some constant k. Here it varies with twice x, so we have The given point allows us to calculate k: y = k/(2x) 6 = k/(2 * 3) k = 36 The function is y = 36/(2x) y = 18/x When x = 9, y = 18/9 = 2 Answeer is C.

OpenStudy (anonymous):

yess :)

OpenStudy (anonymous):

i asked another

OpenStudy (ivettef365):

sure